OpenStudy (anonymous):

The range of the function f(x)=x+5 is (7,9). What is the function's domain? Is it (2,4) (-2,-4) (12,14) (-12,-14) (0,5)

OpenStudy (lynfran):

the range is the y value we can rewrite the given function f(x)=x+5 to y=x+5 and our given range are y=7 and y=9 so sub in ur y values into the equation to find the x value which are the domain... y=x+5

OpenStudy (lynfran):

7=x+5 and 9=x+5 7-5=x and 9-5=x ? =x and ? =x